UT 08-Nov-2019

  • roll over of non-ARGOS eng tasks not completed from UT 07-Nov-2019
  • highest priority:
    • AGW-1 tests - 1 hour
    • LUCI speckle imaging - 1 hour
    • Test PCS patch for IT 7640 (problems with LUCI/SOUL repoint=yes acquisitions) can be seeing up to 2 arcsec
      • send ACE_ACTIVE preset, do acquireref with repoint=yes; PCS patch will be activated in TCS 2019A
    • ESM tests - 45 min per script

SOUL science test script LUCI-1

UT 06:28 - SOUL testing of Mrk 1014 - R=15.5 AO Ref on-axis (galaxy is AO Ref star)
  • Guider shows 1".2
  • UT 06:41 - still waiting to close loop - thin cirrus passing through
  • UT 06:46 - finally did offset after skipping to 2nd line two times and clicking GO
  • -waiiting for collimation no longer showing on telsvc
  • UT 06:48 -resending preset, not optimizing gains
  • LUCI started taking data - 0089 do not use - bad
  • guider reporting 1" (corrections from AO helping - DIMM 1.2")
  • UT 07:05 - DIMM ~1.2"
  • L1 0089-0101 - K-band Mrk 1014
  • UT 07:12 - finished script

ARGOS Eng

UT 11:20 - running M67 ARGOS (SX) vs. ESM (DX) comparison -

L1 - 0120-0129 - K-band ; 0130-0139 H-band ; 0140-0141 J-band ; K-band 0142-0147- skipped the instrument reconfig and did more K-band

L2 - 0003-0012 - K-band ; 0013-0022 H-band ; 0023-0024 J-band ; K-band 0025-0030 - skipped the instrument reconfig and did more K-band

UT 12:15 - shell RIP'd both sides

ESM-DX recovered - lost elevation information on both sides

UT 12:18 - re-sending science preset to both sides - will skip to J-band after acquisition is complete

UT 12:33 - just recovered from a network slowdown - both AdSecs crashed (not ARGOS or ESM related) - Repository froze up - see night log from OSA.

-unusual note, I had paused the LUCI script and skipped to the next J-band dither position, as soon as both OKs were clicked on the AO side, the LUCI script started without my prompting or my clicking OK.

-going to continue collecting data until we can no longer propagate

UT 12:49 - configuring for J-band - will try to take a few exposures (just noticed it hadn't configured to J-band but kept K-band from preset)

L1 0148-0155 J-band

L2 0031-0038 J-band

UT 13:00 - LUCI seemed to hang after an offset - didn't start data collection - had to skip to item and go

UT 13:11 - finished up - closing enclosure - end of laser operations

-Going to close up then take N30 imaging flats K,H,J (LUCI-1 only)

UT 13:22 - starting cals

L1 0156-0165 - K-band

L1 0166-0175 - H-band

L1 0176-185 - J-band

UT 13:41 - putting LUCIs into safe config - done for tonight

-- BarryRothberg - 07 Nov 2019
Topic revision: r5 - 08 Nov 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback